OpenStudy (anonymous):

A person pushes against a 9.2-kg box with a force of 11.5 N. The frictional force on the box is 2.1 N. What is the acceleration of the box? (1 point) 0.86 m/s2 0.92 m/s2 1.0 m/s2 1.3 m/s2 im seriously not getting this lesson :/

OpenStudy (anonymous):

the box has a forward force of 11.5-2.1= 9.4N A=F/M = 9.4/9.1 = 1.03 m/s2 I believe that's how you should calculate it... not 100% sure about the last part since I haven't done this in a while...

OpenStudy (anonymous):

Ok it seems to make sense to me :)

OpenStudy (anonymous):

Newton's Second Law states that the net force acting on an object is equal to the mass times acceleration. \[\sum F = ma\] Let's draw a free body diagram. \(F_a\) is te applied force from the person, \(F_f\) is the frictional force, positive x is to the right. \(F_a\) is positive since it acts to the right, while \(F_f\) is negative since it acts in the opposite direction. We can now right Newton's Second Law as\[F_a - F_f = ma\]This can be solved for a with a little algebra.